A coin is tossed 72 times. Find the standard deviation for the number of heads that will be tossed.
std = sqrt(npq) = sqrt(72(1/2)(1/2)) = (l/2)(6)sqrt(2) = 3sqrt(2)
------
Cheers,
Stan H.

Answer by stanbon(75887)

A coin is tossed 72 times. Find the standard deviation for the number of heads that will be tossed.
std = sqrt(npq) = sqrt(72(1/2)(1/2)) = (l/2)(6)sqrt(2) = 3sqrt(2)